3. (6) Compute / 30x2y dA where R is the shaded region shown below. JR R -2 -1 L2 -11 S20 p2/2 p 16–22 22 4. (5) Evaluate J-2Please answer both and write out the answer.
Answer #1

30,24 dA They souy dydis al popolo Dia dos - isip y?)ds = $1522.4 dx = 60 ( +6019-491] = 60/20] te - 60 = föx 16 = 320 squar